OpenStudy (anonymous):

PLEASE HELP!!!! ): x+y+z=6 x=2y z=x+1 Solve the system of equations... The answer should be (2,1,3) but im so confused on this lesson. Can someone PLEASE help!

OpenStudy (mathmale):

You have a system of three equations in three unknowns, x, y and z. Why not keep the variable x and eliminate the other two variables, y and z? If x = 2y, then y = ? You could substitute this result for y in the first equation. z=x+1. substitute this result into the first equation. Then solve the re-worked first equation for x. Next, find y and z.

OpenStudy (texaschic101):

x = 2y z = x + 1 x + y + z = 6 2y + y + x + 1 = 6 3y + x + 1 = 6 3y + x = 5 3y = -x + 5 y = -1/3x + 5/3 x = 2(-1/3x + 5/3) x = -2/3x + 10/3 x + 2/3x = 10/3 3/3x + 2/3 = 10/3 5/3x = 10/3 x = (10/3) / (5/3) x = 10/3 * 3/5 x = 30/15 x = 2 x = 2y 2 = 2y 2/2 = y 1 = y z = x + 1 z = 2 + 1 z = 3

OpenStudy (texaschic101):

mathmale's way sounds easier

OpenStudy (anonymous):

i get it now! thanks!!! i was having trouble knowing which equation to start with and how to proceed!

OpenStudy (anonymous):

Can you help me with another section??